# Tackling Cyberattacks through AI-based Reactive Systems: A Holistic Review and Future Vision

### v1

@@ VII Conclusions and future work
Human dependence on the use of information technologies has grown exponentially. This dependence poses considerable security challenges, which can be seen in frequent cyberattacks targeting businesses, critical infrastructures, and companies, and even impacting geopolitical events. In response, both private enterprises and nation-states are actively exploring effective ways to strengthen the security measures of their systems, aiming to safeguard sensitive data and prevent unauthorized access. This trend has triggered an increase in research efforts to address these critical challenges. The emergence of cutting-edge technologies, like AI, has enriched the landscape. When integrated into the cybersecurity sector, this technology offers numerous advantages, such as the identification of threats and the identification of appropriate countermeasures. With the assistance of AI, dedicated systems can be used to detect, prevent, and respond to threats within the realm of cyberspace.
@@
Acknowledgement
This work has been partially funded by the strategic project CDL-TALENTUM from the Spanish National Institute of Cybersecurity (INCIBE) and by the Recovery, Transformation and Resilience Plan, Next Generation EU.

### v2

@@ VII Conclusions and future work
Human dependence on the use of information technologies has grown exponentially. This dependence poses considerable security challenges, which can be seen in frequent cyber-attacks targeting businesses, critical infrastructures, and companies, and even impacting geopolitical events. In response, both private enterprises and nation-states are actively exploring effective ways to strengthen the security measures of their systems, aiming to safeguard sensitive data and prevent unauthorized access. This trend has triggered an increase in research efforts to address these critical challenges. The emergence of cutting-edge technologies, like AI, has enriched the landscape. When integrated into the cybersecurity sector, this technology offers numerous advantages, such as the identification of threats and the identification of appropriate countermeasures. With the assistance of AI, dedicated systems can be used to detect, prevent, and respond to threats within the realm of cyberspace.
@@
However, the survey has also identified certain gaps in the research. One of the most notable is the diversity of methodologies and algorithms presented in each paper. This lack of standardization poses a major challenge, as it makes it difficult to compare the different papers analyzed.
From a practical point of view, this lack of uniformity can complicate interoperability. In scenarios such as the creation of a federated CERT (Computer Emergency Response Team), where collaboration and information sharing are essential, the diversity of methodologies and algorithms could obstruct the integration and sharing of data between the different entities. To address this challenge, more attention is needed on standardization and research collaboration. Adopting common standards and encouraging collaboration between researchers can help establish best practices and facilitate comparison and knowledge sharing in the field of cyber threat reaction.
@@
Acknowledgment
This work has been partially funded by the strategic project CDL-TALENTUM from the Spanish National Institute of Cybersecurity (INCIBE) and by the Recovery, Transformation and Resilience Plan, Next Generation EU, and from the postdoctoral grant Margarita Salas (172/MSJD/22).